OpenStudy (anonymous):

solve each system using substitution .. y-2x=3 and 3x-2y=5

OpenStudy (anonymous):

y=2x+3 so 2y=4x+6 and 3x-2y=3x-4x-6=5 and -x=11 so x=-11

OpenStudy (anonymous):

thanks(: